### v2

# What Does Stable Diffusion Know about the 3D Scene?

Abstract
Recent advances in generative models like Stable Diffusion (Rombach et al., 2022) enable the generation of highly photo-realistic images. Our objective in this paper is to probe the diffusion network to determine to what extent it ‘understands’ different properties of the 3D scene depicted in an image. To this end, we make the following contributions: (i) We introduce a protocol to evaluate whether features of an off-the-shelf diffusion model encode a number of physical ‘properties’ of the 3D scene, by training discriminative classifiers on the features for these properties. The probes are applied on datasets of real images with annotations for the property. (ii) We apply this protocol to properties covering scene geometry, scene material, support relations, lighting, and view dependent measures. (iii) We find that features from Stable Diffusion are good for discriminative learning of a number of properties, including scene geometry, support relations, shadows and depth, but less performant for occlusion and material. (iv) We also apply the probes to other networks trained at large-scale, including DINO, CLIP and VQGAN, and find that DINOv2 has a similar performance to Stable Diffusion, while outperforming DINOv1, CLIP and VQGAN.
3D Physical Scene Understanding, Stable Diffusion, Representation Learning
1 Introduction
Figure 1: Motivation: What does Stable Diffusion know about the 3D scene? The Stable Diffusion inpainting model is here tasked with inpainting the masked region of the real images. It correctly predicts a shadow consistent with the lighting direction (top), and a supporting structure consistent with the scene geometry (bottom). This indicates that the Stable Diffusion model generation is consistent with the geometry (of the light source direction) and physical (support) properties. Note: these examples are only for illustration and inpainting is not the objective of this paper. Rather, we probe a general Stable Diffusion network to determine whether there are explicit features for such 3D scene properties. The appendix provides more examples of Stable Diffusion’s capability to predict different physical properties of the scene.
Image generation with diffusion models (Sohl-Dickstein et al., 2015), following on from earlier generation using GANs (Goodfellow et al., 2014), has achieved amazing results in terms of verisimilitude (Rombach et al., 2022). This naturally raises the question of to what extent the diffusion network is modelling the 3D scene depicted in the image. Our objective in this paper is to investigate this question, and we do this precisely by determining whether features from the diffusion network can be used to predict the true geometric and physical properties of the 3D scene. If they can, then that is evidence that the 3D scene is correctly modelled. As an indication that the diffusion network is 3D and physics aware, Figure 1 shows the result of the off-the-shelf Stable Diffusion model (Rombach et al., 2022) inpainting masked regions in real images – it correctly predicts shadows and supporting structures.
To answer this question, we propose an evaluation protocol to systematically ‘probe’ a diffusion network on its ability to represent a number of ‘properties’ of the 3D scene and viewpoint. These properties include: 3D structure and material of the scene, such as surface layout; lighting, such as object-shadow relationships; and viewpoint dependent relations such as occlusion and depth.
The protocol involves three steps: First, a suitable real image evaluation dataset is selected that contains ground truth annotations for the property of interest, for example the SOBA dataset (Wang et al., 2020) is used to probe the understanding of lighting, as it has annotations for object-shadow associations. This provides a train/val/test set for that property; Second, a grid search is carried out over the layers and time steps of the diffusion model to select the optimal feature for determining that property. The selection involves learning the weights of a simple linear classifier for that property (e.g. ‘are these two regions in an object-shadow relationship or not’); Third, the selected feature (layer, time step) and trained classifier are evaluated on a test set, and its performance answers the question of how well the diffusion model ‘understands’ that property.
In short, we probe scene geometry, material, support relation, shadow, occlusion and depth, to answer the question “What does Stable Diffusion Know about the 3D scene?”. We apply this protocol not only to features from the Stable Diffusion network, but also to features from other networks that are trained at large scale, including OpenCLIP (Radford et al., 2021; Ilharco et al., 2021), DINOv1 (Caron et al., 2021), DINOv2 (Oquab et al., 2023), and VQGAN (Esser et al., 2021). This covers networks trained generatively (Stable Diffusion), with self-supervision (DINOv1 & DINOv2), with weak supervision (OpenCLIP), and by auto-regression (VQGAN).
From our investigation, we make two observations: First, the Stable Diffusion and DINOv2 networks have a good ‘understanding’ of the scene geometry, support relations, the lighting, and the depth of a scene, with Stable Diffusion and DINOv2 having a similar and high prediction performance for these properties. However, their prediction of material and occlusion is poorer. Second, Stable Diffusion and DINOv2 generally demonstrate better performance for 3D properties, than other networks trained at large scale: OpenCLIP, DINOv1, and VQGAN.
Why is an understanding of the networks’ ability to predict 3D properties useful? There are four reasons: (1) It begins to answer the scientific question of the extent to which these networks implicitly model the 3D scene; (2) The features we determine that are able to predict 3D physical properties can be used for this task, e.g. to predict shadow-object associations or support relations. This could either be carried out directly by incorporating them in a prediction network, in the manner of Zhan et al. (2023); or they can be used indirectly as a means to train a feed forward network to predict the properties (Wu et al., 2023a, b); (3) By knowing what properties Stable Diffusion is not good at, we have a way to spot images generated by Stable Diffusion, as has been done by Sarkar et al. (2023); (4) It also reveals which properties the network could be trained further on to improve its 3D modelling, e.g., via extra supervision for that property.
We describe the properties explored, the protocol, datasets and classifiers in Section 3. Experimental results of both Stable Diffusion and other networks are given in Section 4, and finally future work is discussed in Section 5.
Table 1: Overview of the datasets and training/evaluation statistics for the properties investigated. For each property, we list the image dataset used, and the number of images for the train, val, and test set. 1000 images are used for testing if the original test set is larger than 1000 images. Regions are selected in each image, and pairs of regions are used for all the probe questions.
Property: Same Plane Perpendicular Plane Material Support Relation Shadow Occlusion Depth
Dataset: ScanNetv2 ScanNetv2 DMS NYUv2 SOBA Sep. COCO NYUv2
Images # Train 400 400 400 400 400 400 400
# Val 100 100 100 100 100 100 100
# Test 1000 1000 1000 654 160 983 654
Regions # Train 7600 4493 4997 8943 3576 6799 8829
# Val 1844 1112 1180 1968 976 1677 2075
# Test 17159 10102 11364 13968 1176 16993 13992
Pairs # Train 14360 17530 18520 13992 7152 19238 15322
# Val 3498 4232 4284 2874 1952 4724 3786
# Test 32654 38640 41824 21768 2352 44266 24880
2 Related Work
2.1 Generative Models
Generative models have made significant achievements in advancing image quality and diversity in the recent literature. A series of generative models, such as Variational Autoencoders (VAEs) (Kingma & Welling, 2014), Generative Adversarial Networks (GANs) (Goodfellow et al., 2014), Flow-based Generators (Dinh et al., 2014), and Diffusion Probabilistic Models (DPMs) (Sohl-Dickstein et al., 2015), have been proposed. These models have contributed to widespread tasks, including image completion (Pathak et al., 2016), composition (Lin et al., 2018), interpolation (Karras et al., 2019) and editing (Chai et al., 2021), image-to-image translation (Isola et al., 2017), multi-modalities translation (Hu et al., 2023), and numerous others. We build upon the diffusion models (Rombach et al., 2022), which have shown state-of-the-art generation quality.
2.2 Exploration of Pre-trained Models
Building on the success of deep networks, there has been significant interest from the community to understand what has been learnt by these complex models. On discriminative models, for example, Zeiler & Fergus (2014); Mahendran & Vedaldi (2015) propose inverse reconstruction to directly visualize the acquired semantic information in various layers of a trained classification network; Zhou et al. (2016); Fong & Vedaldi (2017); Fong et al. (2019) demonstrate that scene classification networks have remarkable localization ability despite being trained on only image-level labels; and Erhan et al. (2009); Simonyan et al. (2014); Selvaraju et al. (2017) use the gradients of any target concept, flowing into the final convolutional layer to produce a saliency map highlighting important regions in the image for predicting the concept. In the more recent literature, Chefer et al. (2021) explore what has been learned in the powerful transformer model by visualizing the attention map. On generative models, researchers have mainly investigated what has been learned in GANs, for example, GAN dissection (Bau et al., 2019) presents an analytic framework to visualize and understand GANs at the unit-, object-, and scene-level; Wu et al. (2021) analyse the latent style space of StyleGANs (Karras et al., 2019). The most recent work (Sarkar et al., 2023) studies the 3D geometric relations in generated images, such as vanishing points and shadows, and notes that the errors made can be used to discriminate real from generated images.
2.3 Exploitation of Generative Models
Apart from understanding the representation in pre-trained models, there has been a recent trend for exploiting the learnt feature from generative models, to tackle a series of downstream discriminative tasks. For example, leveraging generative models for data augmentation in recognition tasks (Jahanian et al., 2022; He et al., 2023), semantic segmentation via generative models (Li et al., 2021; Baranchuk et al., 2022; Xu et al., 2023), open-vocabulary segmentation with diffusion models (Li et al., 2023), depth maps estimation based on RGB images (Noguchi & Harada, 2020; Shi et al., 2022). More recently, Bhattad et al. (2023) search for intrinsic offsets in a pre-trained StyleGAN for a range of downstream tasks, predicting normal maps, depth maps, segmentations, albedo maps, and shading. In contrast to this work, we adopt annotations from different datasets for supervision, rather than employing pre-trained prediction models for supervision. A closely related effort to ours is DIffusion FeaTures (DIFT) (Tang et al., 2023), but it only focuses on computing correspondences at the geometric or semantic level between images.
2.4 Physical Scene Understanding
There have been works studying different physical properties for scene understanding, including shadows (Wang et al., 2020, 2021), material (Upchurch & Niu, 2022; Sharma et al., 2023), occlusion (Zhan et al., 2022), scene geometry (Liu et al., 2019), support relations (Silberman et al., 2012) and depth (Silberman et al., 2012). However, these works focus on one or two physical properties, and most of them require training a model for the property in a supervised manner. In contrast, we use a single model to predict multiple properties, and do not train the features.
3 Method – Properties, Datasets, and Classifiers
Figure 2: Example images for probing scene geometry. The first row shows a sample annotation for the same plane, and the second row is a sample annotation for perpendicular plane. Here, and in the following figures, (A, B) are a positive pair, while (A, C) are negative. The images are from the ScanNetv2 dataset (Dai et al., 2017) with annotations for planes from (Liu et al., 2019). In the first row, the first piece of floor (A) is on the same plane as the second piece of floor (B), but is not on the same plane as the surface of the drawers (C). In the second row, the table top (A) is perpendicular to the wall (B), but is not perpendicular to the stool top (C).
Figure 3: Example images for probing material, support relation and shadow. The first row is for material, the second row for support relation, and the third row for shadow. First row: the material images are from the DMS dataset (Upchurch & Niu, 2022). The paintings are both covered with glass (A and B) whereas the curtain (C) is made of fabric. Second row: the support relation images are from the NYUv2 dataset (Silberman et al., 2012). The paper (A) is supported by the table (B), but it is not supported by the chair (C). Third row: the shadow images are from the SOBA dataset (Wang et al., 2020). The person (A) has the shadow (B), not the shadow (C).
Our goal is to examine the ability of a diffusion model to understand different physical properties of the 3D scene, including: scene geometry, material, support relations, shadows, occlusion and depth. Specifically, we conduct linear probing of the features from different layers and time steps of the Stable Diffusion model. First, we set up the questions for each property (Section 3.1); and then select real image datasets with ground truth annotations for each property (Section 3.2). We describe how a classifier is trained to answer the questions, and the grid search for the optimal time step and layer to extract a feature for predicting the property in Section 3.3.
3.1 Properties and Questions
Here, we study the diffusion model’s ability to predict different properties of the 3D scene; the properties cover the 3D structure and material, the lighting, and the viewpoint. For each property, we propose questions that classify the relationship between a pair of Regions, A and B, in the same image, based on the features extracted from the diffusion model. The properties and questions are:
  • Same Plane: ‘Are Region A and Region B on the same plane?’
  • Perpendicular Plane: ‘Are Region A and Region B on perpendicular planes?’
  • Material: ‘Are Region A and Region B made of the same material?’
  • Support Relation: ‘Is Region A (object A) supported by Region B (object B)?’
  • Shadow: ‘Are Region A and Region B in an object-shadow relationship?’
  • Occlusion: ‘Are Region A and Region B part of the same object but, separated by occlusion?’
  • Depth: ‘Does Region A have a greater average depth than Region B?’
We choose these properties as they exemplify important aspects of the 3D physical scene: the Same Plane and Perpendicular Plane questions probe the 3D scene geometry; the Material question probes what the surface is made of, e.g., metal, wood, glass, or fabric, rather than its shape; the Support Relation question probes the physics of the forces in the 3D scene; the Shadow question probes the lighting of the scene; the Occlusion and Depth questions depend on the viewpoint, and probe the disentanglement of the 3D scene from its viewpoint.
3.2 Datasets
To study the different properties, we adopt various off-the-shelf real image datasets with annotations for the different properties, where the dataset used depends on the property. We repurpose each dataset to support probe questions of the form: $\mathcal{D}=\{(R_{A},R_{B},y)_{1},\dots,(R_{A},R_{B},y)_{n}\}$, where $R_{A}$, $R_{B}$ denote a pair of regions, and $y$ is the binary label indicating the answer to the considered question of the probed property. For each property, we create a train/val/test split from those of the original datasets, if all three splits are available. While for dataset with only train/test splits available, we divide the original train split into our train/val splits. Table 1 summarises the datasets used and the statistics of the splits used. We discuss each property and dataset in more detail next.
Same Plane. We use the ScanNetv2 dataset (Dai et al., 2017) with annotations for planes from Liu et al. (2019). Regions are obtained via splitting plane masks into several regions. A pair of regions are positive if they are on the same plane, and negative if they are on different planes. First row of Figure 2 is an example.
Perpendicular Plane. We use the ScanNetv2 dataset (Dai et al., 2017). We use the annotations from Liu et al. (2019) which provide segmentation masks as well as plane parameters for planes in the image, so that we can obtain the normal of planes to judge whether they are perpendicular or not. A pair of regions are positive if they are on perpendicular planes, and negative if they are not on perpendicular planes. Second row of Figure 2 is an example.
Material. We adopt the recent DMS dataset (Upchurch & Niu, 2022) to study the material property, which provides dense annotations of material category for each pixel in the images. Therefore, we can get the mask of each material via grouping pixels with the same material label together. In total, there are 46 pre-defined material categories. Regions are obtained by splitting the mask of each material into different connected components, i.e., they are simply groups with same material labels, yet not connected. A pair of regions are positive if they are of the same material category, and negative if they are of different material categories. First row of Figure 3 is an example.
Support Relation. We use the NYUv2 dataset (Silberman et al., 2012) to probe the support relation. Segmentation annotations for different regions (objects or surfaces) are provided, as well as their support relations. Support relation here means an object is physically supported by another object, i.e., the second object will undertake the force to enable the first object to stay at its position. Regions are directly obtained via the segmentation annotations. A pair of regions are positive if the first region is supported by the second region, and negative if the first region is not supported by the second region. Second row of Figure 3 is an example.
Shadow. We use the SOBA dataset (Wang et al., 2020, 2021) to study the shadows which depend on the lighting of the scene. Segmentation masks for each object and shadow, as well as their associations are provided in the dataset annotations. Regions are directly obtained from the annotated object and shadow masks. In a region pair, there is one object mask and one shadow mask. A pair of regions are positive if the shadow mask is the shadow of the object, and negative if the shadow mask is the shadow of another object. Third row of Figure 3 is an example.
Figure 4: Example images for probing viewpoint-dependent properties (occlusion & depth). The first row is for occlusion and the second row is for depth. First row: the occlusion images are from the Separated COCO dataset (Zhan et al., 2022). The sofa (A) and the sofa (B) are part of the same object, whilst the monitor (C) is not part of the sofa. Second row: the depth images are from the NYUv2 dataset (Silberman et al., 2012). The chair (A) is farther away than the object on the floor (B), but it is closer than the cupboard (C).
Occlusion. We use the Seperated COCO dataset (Zhan et al., 2022) to study the occlusion (object seperation) problem. Regions are different connected components of objects (and the object mask if it is not separated), i.e., groups of connected pixels belonging to the same object. A pair of regions are positive if they are different components of the same object separated due to occlusion, and negative if they are not from the same object. First row of Figure 4 is an example.
Depth. We use the NYUv2 dataset (Silberman et al., 2012), that provides mask annotations for different objects and regions, together with depth for each pixel. A pair of regions are positive if the first region has a greater average depth than the second region, and negative if the first region has a less average depth than the second region. The average depth of a region is calculated via the average of depth value of each pixel the region contains. Second row of Figure 4 is an example.
3.3 Property Probing
We aim to determine which Stable Diffusion features best represent different properties. To obtain features from an off-the-shelf Stable Diffusion network, we follow the approach of Tang et al. (2023) used for DIFT, where noise is added to the input image in the latent space, and features are extracted from different layers and time steps of the model. While probing the properties, linear classifiers are used to infer the relationships between regions. The region representation is obtained by a simple average pooling of the diffusion features over the annotated region or object.
Extracting Stable Diffusion Features. We add noise $\epsilon\sim\mathcal{N}(0,\mathbf{I})$ of time step $t\in[0,T]$ to the input image $x_{0}$’s latent representation $z_{0}$ encoded by the VAE encoder:
$z_{t}=\sqrt{\alpha_{t}}z_{0}+(\sqrt{1-\alpha_{t}})\epsilon$ (1)
and then extract features from the immediate layers of a pre-trained diffusion model, $f_{\theta}(\cdot)$ after feeding $z_{t}$ and $t$ in $f_{\theta}$ ($f_{\theta}$ is a U-Net consisting of 4 downsampling layers and 4 upsampling layers):
$F_{t,l}=f_{\theta_{l}}(z_{t},t)$ (2)
where $f_{\theta_{l}}$ is the $l$-th U-Net layer. In this way, we can get the representation of an image $F_{t,l}$ at time step $t$ and $l$-th U-Net layer for the probe. We upsample the obtained representation to the size of original image with bi-linear, then use the region mask to get a region-wise feature vector, by averaging the feature vectors of each pixel it contains, i.e., average pooling.
$v_{k,t,l}=\text{avgpool}(R_{k}\odot\text{upsample}(F_{t,l}))$ (3)
where $v_{k,t,l}$ is the feature vector of the $k$-th region $R_{k}$. $\odot$ here is a per-pixel product of the region mask and the feature.
Figure 5: (a) Nomenclature for the U-Net Layers. We probe 4 downsampling encoder layers $E_{1}$-$E_{4}$ and 4 upsampling decoder layers $D_{1}$-$D_{4}$ of the Stable Diffusion U-Net. (b) A prediction failure for Material. In this example the model does not predict that the two regions are made of the same material (fabric). (c) A prediction failure for Occlusion. In this example the model does not predict that the two regions belong to the same object (the sofa).
Linear Probing. After computing features from a diffusion model, we use a linear classifier (a linear SVM) to examine how well these features can be used to answer questions to each of the properties. Specifically, the input of the classifier is the difference or absolute difference between the feature vectors of Region A and Region B, i.e., $v_{A}-v_{B}$ or $|v_{A}-v_{B}|$, and the output is a Yes/No answer to the question. Denoting the answer to the question as $Q$, then since the questions about Same Plane, Perpendicular Plane, Material, Shadow and Occlusion are symmetric relations, $Q(v_{A},v_{B})=Q(v_{B},v_{A})$. However, the questions about Support Relation and Depth are not symmetric. Thus, we use $|v_{A}-v_{B}|$ (a symmetric function) as input for the first group of questions, and $v_{A}-v_{B}$ (non-symmetric) for the rest of questions. We train the linear classifier on the train set via the positive/negative samples of region pairs for each property; do a grid search on the validation set to find (i) the optimal time step $t$, (ii) the U-Net layer $l$, and (iii) the SVM regularization parameter $C$; and evaluate the performance on the test set.
Discussion. Some of the current symmetric questions can be reformulated in a non-symmetric manner in order to obtain more information about the property. For example, the shadow question could be formulated as ‘is region A the shadow of object B’ rather than the (symmetric) ‘are region A and region B in an object-shadow relationship’. The non-symmetric version requires the classifier to explicitly identify which region is the object, and which the shadow. Note, the protocol has been explained for diffusion models, but in the following we also apply it to other networks trained at large-scale. In Section 4.3 we give results for its application to OpenCLIP (Radford et al., 2021; Ilharco et al., 2021), DINOv1 (Caron et al., 2021), DINOv2 (Oquab et al., 2023), and VQGAN (Esser et al., 2021).
4 Experiments
In this section, we first give details of the grid search method in Section 4.1. We then give the linear probing grid search results on features from Stable Diffusion in Section 4.2 and from other networks trained at scale in Section 4.3. Finally, we compare all models on the test set in Section 4.4.
4.1 Implementation Detail and Evaluation Metric
Implementation Details. For each property, we sample the same number of positive / negative pairs, to maintain a balanced evaluation set. In terms of the linear SVM, we tune the penalty parameter $C$ on the val split to find the best $C$ for each property. Therefore, we are grid searching 3 parameters on the val set, namely, Stable Diffusion Timestep $t$ ranging from 1 to 1000, U-Net Layer $l$ covering the 4 downsampling and 4 upsampling layers, and the SVM penalty parameter $C$ ranging over $0.001,0.01,0.1,1,10,100,1000$. The timestep is searched with a stride of 20 steps, since the difference in performance around the optimal value varies by less than 0.01 within 20 steps. In practice the $C$ parameter is always between $0.1$ and $1$, so we carry out a finer search over values between $0.1$ and $1.0$ in steps of $0.1$. The linear SVM is solved using the libsvm library (Chang & Lin, 2011) with the SMO algorithm, to get the unique global optimal solution. Please refer to the appendix for more implementation details.
Evaluation Metric. All protocols are binary classification, therefore, we use ROC Area Under the Curve (AUC Score) to evaluate the performance of the linear classifier, as it is not sensitive to different decision thresholds.
Table 2: SVM grid search results of Stable Diffusion features. For each property, we train the linear SVM on the training set and grid search the best combination of time step, layer, and $C$ on the validation set. The ROC AUC score (%) is reported on the validation set using the selected combination.
Property Time Step Layer $C$ Val AUC
Same Plane 360 $D_{3}$ 0.4 97.3
Perpendicular Plane 160 $D_{3}$ 0.5 88.5
Material 20 $D_{3}$ 0.5 81.5
Support Relation 120 $D_{3}$ 1.0 92.6
Shadow 160 $D_{3}$ 0.8 95.4
Occlusion 180 $D_{3}$ 0.3 83.8
Depth 60 $D_{3}$ 0.9 99.2
Table 3: SVM grid search results of other features trained at scale. We train the linear SVM on the training set, and grid search the best combination of ViT/Transformer layer and $C$ on the validation set. The OpenCLIP and VQGAN models we use have 48 transformer layers, DINOv1 has 12 layers and DINOv2 has 40 layers. The $i$-th layer means the $i$-th transformer layer from the input side.
Material Support Relation
OpenCLIP DINOv1 DINOv2 VQGAN OpenCLIP DINOv1 DINOv2 VQGAN
Optimal Layer 30 8 23 11 32 9 40 14
Optimal C 0.3 0.2 0.6 0.3 0.3 0.3 0.6 0.4
Val AUC 77.5 77.4 81.3 65.8 92.0 91.5 93.6 85.4
Table 4: Performance of Stable Diffusion features compared to state-of-the-art self-supervised features. For each property, we use the best time step, layer and $C$ found in the grid search in Table 2 for Stable Diffusion, and the best layer and $C$ found in the grid search for other features. The performance is the ROC AUC on the test set, and ‘Random’ means a random classifier.
Property Random OpenCLIP DINOv1 DINOv2 VQGAN Stable Diffusion
Same Plane 50 92.3 91.4 94.5 81.3 96.3
Perpendicular Plane 50 71.8 71.3 82.1 62.0 86.0
Material 50 78.7 78.8 83.5 65.5 83.6
Support Relation 50 90.6 90.8 92.8 84.1 92.1
Shadow 50 94.9 92.2 95.8 89.0 95.4
Occlusion 50 81.2 79.9 84.4 78.4 84.8
Depth 50 99.2 97.1 99.7 91.8 99.6
4.2 Results for Stable Diffusion
The results for grid search are shown in Table 2. For Stable Diffusion U-Net Layer, $D_{l}$ means the $l$-th layer of the U-Net decoder, i.e., upsampling layer, from outside to inside (right to left), and we provide an illustration of the layers in Figure 5(a). We can make two observations: First, the best time step for different properties varies, while the best U-Net layer is always $D_{3}$ in the decoder rather than the encoder. Further explorations using Stable Diffusion features for downstream tasks could thus start from the U-Net decoder layers, especially $D_{3}$. Second, in terms of the performance on the test set, we find that Stable Diffusion can understand very well about scene geometry, support relations, shadows, and depth, but it is less performant at predicting material and occlusion. Examples of its failure are shown in Figure 5 (b) (c). As noted in Zhan et al. (2022) and Kirillov et al. (2023), grouping all separated parts of an object due to occlusion remains challenging even for state-of-the-art detection and segmentation models.
4.3 Results for Other Features Trained at Scale
We have also applied our protocol to other models pre-trained on large scale image datasets, including OpenCLIP (Radford et al., 2021; Ilharco et al., 2021) pre-trained on LAION dataset (Schuhmann et al., 2022), DINOv1 (Caron et al., 2021) pre-trained on ImageNet dataset (Deng et al., 2009), DINOv2 (Oquab et al., 2023) pre-trained on LVD-142M dataset (Oquab et al., 2023), and VQGAN (Esser et al., 2021) pre-trained on ImageNet dataset (Deng et al., 2009). We use the best pre-trained checkpoints available on their official GitHub – ViT-B for DINOv1, ViT-G for OpenCLIP and DINOv2, and the 48-layer transformer checkpoint for VQGAN. Similar to Stable Diffusion, for each of these models, we conduct a grid search on the validation set in terms of the ViT/Transformer layer and $C$ for SVM, and use the best combination of parameters for evaluation on the test set. Grid search results are reported in Table 3 for the tasks of material and support. Results of other tasks are in the appendix.
4.4 Comparison of Different Features Trained at Scale
We compare the state-of-the-art pre-trained large-scale models’ representations on various downstream tasks in Table 4. It can be observed that the Stable Diffusion and DINOv2 representations outperform OpenCLIP, DINOv1 and VQGAN for all properties, indicating the potential of utilizing Stable Diffusion and DINOv2 representations for different downstream tasks with the optimal time steps and layers found in Section 4.2 and Section 4.3.
5 Discussion and Future Work
In this paper, we have developed a protocol to examine whether the Stable Diffusion and other models pre-trained on large scale image datasets, like CLIP, DINO and VQGAN, have explicit feature representations for different properties of the 3D scene.
It is interesting to find that different time steps of Stable Diffusion and different layers of DINOv2 representations can handle several different physical properties at a state-of-the-art performance, indicating the potential of utilising the Stable Diffusion and DINOv2 models for different downstream tasks. However, for some properties such as material and occlusion, the networks are not distilling the information in a manner that can be used by a linear probe. This could indicate that these properties are not modelled well by the network or that more than a linear probe is required to tease them out. We show examples of the prediction failures for these properties in Figure 5. In the appendix, we show that such prediction failures also occur in generated (i.e. synthetic) images. It is worth noting that occlusion is a challenge even for the powerful Segment Anything Model (SAM) (Kirillov et al., 2023), where the model ‘hallucinates small disconnected components at times’.
This paper has given some insight into answering the question: ‘What does Stable Diffusion know about the 3D scene?’ for real images. Of course, there are more properties that could be investigated in the manner proposed here. For example, contact relations (Fouhey et al., 2016) and object orientation (Xiang et al., 2018), as well as the more nuanced non-symmetric formulations of the current questions. Another direction would be to use the pixel-wise supervision method of Bhattad et al. (2023) to search for features that can predict depth maps, normal maps, etc. Furthermore, utilising combinations of features from different time steps and layers as in Luo et al. (2023); Zhang et al. (2023) could be explored to obtain even better performance for using Stable Diffusion features for downstream tasks.
Acknowledgements.
[…]
Appendix
Appendix A More Implementation Details
Extracting Stable Diffusion Features.
Following DIFT (Tang et al., 2023), when we extract Stable Diffusion features, we add a different random noise 8 times and then take the average of the generated features. The process can be completed in one forward pass as we are using a batch of 8. We use an empty prompt ‘’ as the text prompt.
Train/Val Partition.
For the partition of train/val split, we select the train & val images from different scenes for the NYUv2 (Silberman et al., 2012) and ScanNetv2 (Dai et al., 2017) dataset.
Sampling of Images.
For the train/val/test splits, if the number of images used is less than the original number of images in the datasets, we randomly sample our train/val/test images from the original datasets.
Sampling of Positive/Negative Pairs.
For each image, if the number of possible negative pairs is larger than the number of possible positive pairs, we randomly sample from the negative pairs to obtain an equal number of negative and positive pairs, and vice versa. In this way, we keep a balanced sampling of positive and negative pairs for the binary linear classifier. As can be observed in Table 1, the number of train/val pairs for different properties are different, although we keep the same number of train/val images for different properties. This is because for different properties the availability of positive/negative pairs are different. For depth, we select a pair only if the average depth of one region is 1.2 times greater than the other because it is even challenging for humans to judge the depth order of two regions below this threshold. For perpendicular plane, taking the potential annotation errors into account, we select a pair as perpendicular if the angle between their normal vectors is greater than 85°and smaller than 95°, and select a pair as not perpendicular if the angle between their normal vectors is smaller than 60°or greater than 120°.
Region Filtering.
When selecting the regions, we filter out the small regions, e.g., regions smaller than 100 pixels, because regions that are too small are challenging even for humans to annotate.
Image Filtering.
As there are some noisy annotations in the (Liu et al., 2019) dataset, we manually filter the images whose annotations are inaccurate.
Linear SVM.
The feature vectors are L2-normalised before inputting into the linear SVM. The binary decision of the SVM is given by $sign(w^{T}v+b)$, where $v$ is the input vector to SVM:
$v=|v_{A}-v_{B}|$ (4)
for the Same Plane, Perpendicular Plane, Material, Shadow and Occlusion questions, and
$v=v_{A}-v_{B}$ (5)
for the Support Relation and Depth questions.
Extension of Separated COCO.
To study the occlusion problem, we utilise the Separated COCO dataset (Zhan et al., 2022). The original dataset only collects separated objects due to occlusion in the COCO 2017 val split, we further extend it to the COCO 2017 train split for more data using the same method as in (Zhan et al., 2022). As the original COCO val split is too small (only 5k images v.s. 118k images in the COCO train split), we get our partition of train/val/test splits by dividing the generated dataset ourselves.
Appendix B Analysis of Stable Diffusion Generated Images
As Figure 1 shows, our motivation for the paper is that we observe that Stable Diffusion correctly predicts different physical properties of the scene. The reason why we do not study the generated images directly is that there are no annotations available on different properties for these synthetic images, so it is expensive to get quantitative results. But in this section, we provide more qualitative examples and analysis of Stable Diffusion generated images in terms of different physical properties. The observations match our findings on real images – Stable Diffusion ‘knows’ about a number of physical properties including scene geometry, material, support relations, shadows, occlusion and depth, but may fail in some cases in terms of material and occlusion.
We show examples for: Scene Geometry in Figure 6; Material, Support Relations, and Shadows in Figure 7; and Occlusion and Depth in Figure 8.
Figure 6: Stable Diffusion generated images testing scene geometry prediction. Here and for the following figures, the model is tasked with inpainting the masked region of the real images. Stable Diffusion ‘knows’ about same plane and perpendicular plane relations in the generation. When the intersection of two sofa planes (first row), two walls (second and sixth row), two cabinet planes (third row), two pillar planes (fourth row) or two fridge planes (fifth row) is masked out, Stable Diffusion is able to generate the two perpendicular planes at the corner based on the unmasked parts of the planes.
Figure 7: Stable Diffusion generated images testing material, support relation and shadow prediction. Stable Diffusion ‘knows’ about support relations and shadows in the generation, but may fail sometimes for material. Rows 1-2: Material; Rows 3-4: Support Relation; Rows 5-6: Shadow. In the first row, the model distinguishes the two different materials clearly and there is clear boundary between the generated pancake and plate; while in the second row, the model fails to distinguish the two different materials clearly, generating a mixed boundary. In the third row and fourth rows, the model does inpaint the supporting object for the stuff on the table and the machine. In the fifth and sixth rows, the model manages to inpaint the shadow correctly. Better to zoom in for more details.
Figure 8: Stable Diffusion generated images testing occlusion and depth prediction. Stable Diffusion ‘knows’ about depth in the generation, but may fail sometimes for occlusion. Rows 1-3: Occlusion; Rows 4-6: Depth. In Row 1, the model fails to connect the tail with the cat body and generates a new tail for the cat, while in Row 2, the model successfully connects the separated people and generates their whole body, and in Row 3, the separated parts of oven are connected to generate the entire oven. In Rows 4-6, the model correctly generates a car of the proper size based on depth. The generated car is larger if it is closer, and smaller if it is farther away.
Appendix C Additional Results for Other Features Trained at Large Scale
As mentioned in Section 4.3, we have conducted grid search for all the other large pre-trained models, including OpenCLIP, DINOv1, DINOv2 and VQGAN for all tasks. Tables in this section provide results for the Same Plane (Table 5), Perpendicular Plane (Table 6), Shadow (Table 7), Occlusion (Table 8) and Depth (Table 9) tasks for these models.
Table 5: SVM grid search results of other features trained at large scale for Same Plane task.
Same Plane
OpenCLIP DINOv1 DINOv2 VQGAN
Optimal Layer 27 8 24 12
Optimal C 0.7 0.7 0.8 1.0
Val AUC 94.5 93.2 96.0 82.6
Table 6: SVM grid search results of other features trained at large scale for Perpendicular Plane task.
Perpendicular Plane
OpenCLIP DINOv1 DINOv2 VQGAN
Optimal Layer 27 9 22 12
Optimal C 1.0 0.2 0.6 0.6
Val AUC 72.9 70.9 84.9 62.8
Table 7: SVM grid search results of other features trained at large scale for Shadow task.
Shadow
OpenCLIP DINOv1 DINOv2 VQGAN
Optimal Layer 28 2 29 8
Optimal C 1.0 0.8 1.0 1.0
Val AUC 94.6 92.4 96.6 88.7
Table 8: SVM grid search results of other features trained at large scale for Occlusion task.
Occlusion
OpenCLIP DINOv1 DINOv2 VQGAN
Optimal Layer 31 3 29 2
Optimal C 0.2 0.2 0.3 1.0
Val AUC 80.6 77.0 84.4 77.4
Table 9: SVM grid search results of other features trained at large scale for Depth task.
Depth
OpenCLIP DINOv1 DINOv2 VQGAN
Optimal Layer 32 7 30 45
Optimal C 0.1 0.4 1.0 0.5
Val AUC 99.2 97.4 99.6 93.7
Appendix D Train/Val AUC Results of Stable Diffusion Features
Table 10 shows the train/val AUC of the SVM grid search results for Stable Diffusion features at the best combination of time step, layer and $C$ as in Table 4.2.
Table 10: Train/Val AUC of SVM grid search for Stable Diffusion features. For each property, the Train/Val AUC at the best combination of time step, layer and $C$ is reported.
Property Train AUC Val AUC
Same Plane 97.5 97.3
Perpendicular Plane 90.1 88.5
Material 87.6 81.5
Support Relation 93.7 92.6
Shadow 96.2 95.4
Occlusion 86.7 83.8
Depth 99.8 99.2